OpenStudy (anonymous):

find the limit: lim(x->0) sqrt(x+4)-2/x

OpenStudy (anonymous):

if it is: \(\lim_{x\to 0}\sqrt{x+4}-\frac{2}{x}=-\infty\) if it is: \(\lim_{x\to 0}\frac{\sqrt{x+4}-2}{x}\) use L'Hopital's rule

OpenStudy (accessdenied):

In that second case you might also try this "trick": \( \begin{align} \dfrac{\sqrt{x + 4} - 2}{x} &= \dfrac{\sqrt{x + 4} - 2}{x + 4 - 4} \\ &= \dfrac{\sqrt{x + 4} - 2}{(\sqrt{x + 4} + 2)(\sqrt{x + 4} - 2)} \\ \end{align}\) Which allows you to cancel out the problem value and use direct substitution. :)

OpenStudy (anonymous):

Rationalize the numerator.